Two bodies A and B of equal masses moving with velocities (10km/s, 030°) and (15km/s, 090°) impinges and coalesced. Find the initial velocity with which the composite body begins to move immediately after the collision
1
Expert's answer
2019-10-16T09:18:23-0400

From the conservation of momentum:


"m\bold{v_1}+m\bold{v_2}=2m\bold{v}"

"0.5(10\cos{30},10\sin{30})+0.5(0,15)=(v_x,v_y)"

"v_x=4.33\frac{m}{s}, v_y=10\frac{m}{s}"

Thus,


"v=\sqrt{4.33^2+10^2}=10.9\frac{m}{s}"

"\theta=\arctan{\frac{10}{4.33}}=66.6\degree"